OpenStudy (anonymous):

Solve for x: |x| − 8 = −5

OpenStudy (anonymous):

x>0 then x-8=-5 x=3 x<0 then -x-8=-5 x=-3

OpenStudy (anonymous):

what about this one Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

x>7 then -3(x+7)=-12 x<-7 then 3(x+7)=-12 i think you can solve for x now... (:

OpenStudy (anonymous):

but how do you find out what x is?

OpenStudy (anonymous):

asking about the mod or just algebraic operations?

OpenStudy (anonymous):

in the question

OpenStudy (anonymous):

um asking in which part you are confused?

OpenStudy (anonymous):

how you find out what x is, it could be anything really... and how you get 2 answers for x

OpenStudy (anonymous):

hey if you will graph this equation then you will found two values of x for one y... and that what a mod function is,, i defined mod function as >0 or <0 by substituting x>7 or <7

OpenStudy (anonymous):

okay... why did you choose 7?

OpenStudy (anonymous):

`coz i want to take mod +ive or -ive and neighbours of 7 were capable of it :P

OpenStudy (anonymous):

is -3 the first one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then it would be 11?

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

okay, thanks :D